OpenStudy (anonymous):

The length of one leg of a right triangle is 3x the length of the other. The length of the hypotenuse is 10. What is the length of the longest leg?

OpenStudy (amistre64):

x = x y = 3x x^2 + y^2 = 10^2 x^2 + (3x)^2 = 100 x^2 + 9x^2 = 100 10x^2 = 100 x^2 = 10 x=sqrt(10) y=3sqrt(10) if I did me math right